# Symmetry Breaking in Chemical Systems: Engineering Complexity through Self-Organization and Marangoni Flows

### v3

@@ 2.2 Estimation of dimensionless numbers and timescale comparison
In our experiments, the chemical fingers forming on the pillars or on the walls of the PDMS/PMMA setups exhibit a wavelength of approximately $\lambda\approx 2~{}mm$. We hypothesize that the fluid fingers have a similar wavelength. Assuming the climbing fluid depth is in the range $h=20-100~{}\mu$[77, 78], we calculate the Marangoni stress $\tau$ using formula [69] $h=3\tau/2\rho g$, which represents the balance between the competing convective effects of Marangoni and gravitational forces. This calculation yields a surface stress range of $\tau=0.13-0.67$ $N/m^{2}$. Evaluating a solutal Marangoni number for $\tau=0.13$ $N/m^{2}$ and $h=20~{}\mu m$ yields:
$\mathscr{M}_{S}=\frac{\tau h^{2}}{\mu D}\sim\frac{0.13Nm^{-2}\times(20\times 10^{-6})^{2}m^{2}}{10^{-3}Nsm^{-2}\times 1.2\times 10^{-9}m^{2}s^{-1}}\sim 50,$
where $\mu$ is the viscosity of the CHD-BZ solution and $D$ is molecular diffusivity of chemical components. The thermal Marangoni number, defined as:
@@
is two orders of magnitude smaller due to the ratio of molecular diffusivity to thermal diffusivity, given by
$D/\kappa=(1.2\times 10^{-9}m^{2}s^{-1})/(1.5\times 10^{-7}m^{2}s^{-1})\sim 10^{-2}.$
We now proceed by considering a dimensional analysis similar to Ref. [78] that examines the interaction between a deformable free surface and the steady-state concentration profile, which is governed by diffusion, evaporation, and convective transport along the pillars. Evaporation reduces the concentration of the more volatile chemical component with lower surface tension near the surface; however, under the assumption of a quasi-steady state, fresh fluid from the reservoir continuously replenishes the depleted region, establishing a stable vertical concentration gradient. We note that given the complexity of the CHD-BZ reaction, where multiple chemical components are continuously produced and consumed, the validity of this assumption requires further extensive investigation in future studies. If the free surface experiences an upward perturbation, the concentration profile must adjust before the interface returns to its original shape. If the concentration realigns more rapidly than the interface relaxes, the perturbed region develops a lower concentration than its surroundings, generating a surface tension gradient that draws fluid into the bump and amplifies the perturbation. Similarly, depressions in the free surface create local surface tension minima, pulling fluid away and reinforcing the deformation. This feedback loop sustains convection and enhances surface fluctuations, making the film susceptible to instability driven by Marangoni forces. The system’s stability depends on the relative timescales of diffusion and interface relaxation. If the diffusive timescale is significantly shorter than the relaxation time, the system is unstable, i.e., $T_{\mathrm{diff}}/T_{\mathrm{relax}}\ll 1.$ The relaxation time $T_{\mathrm{relax}}$ is estimated by balancing pressure gradients with viscous stresses,
@@ 2.3 Numerical Simulations
Figure 12: Numerical simulations of the Tyson-Fife model (A) The space-time plot of concentrations $u$, $v$ and $w$ along the line marked in panel (B) illustrates that the obstacles become wave centers. (B) shows an image of the circular waves emanating from the obstacles at dimensionless time $t$$=$$7.5$. Time and space are in arbitrary units. (C) Simulations incorporating an asymmetric distribution of the inhibitor $w$ around the obstacles demonstrate wavefront instability at different time points. Only concentration of activator $u$ is shown (see Figure S6 for concentrations of both slow and fast inhibitors $v$ and $w$).
A full numerical simulation of the system would require solving the Navier-Stokes equations coupled with the chemical reaction dynamics. The Marangoni stress, imposed as a boundary condition on the deformed fluid’s free surface, connects fluid flow to chemical concentrations via surface tension gradients induced by concentration and thermal differences. While such simulations are beyond the scope of this manuscript and are currently being explored in our laboratory, we adopt a simplified approach that omits fluid flow and implicitly accounts for evaporation. Specifically, we assume that the reaction inhibitor (bromide ions) has a lower concentration near the pillars. Bromine gas, a product of the CHD-BZ reaction that mediates bromide ion production, is removed more rapidly in the vicinity of the obstacles, where the fluid ascends the hydrophilic pillars. As a result, the bromide ion concentration is effectively reduced on the pillars compared to the bulk solution. To explore the resulting wave dynamics, we employed the Tyson-Fife model [80] in two dimensions (see Materials and Methods for details). In our simulations, shown in Figure 12A-B, the inhibitor variable $w$ (bromide ions) was set to zero at the locations of the pillars. This localized depletion of the fast inhibitor $w$ near the obstacles leads to an increased concentration of the activator, which in turn facilitates the oxidation of the metal catalyst and the initiation of oxidation waves at the obstacles. These results align with our experimental observations.
We also analyzed the effect of an uneven angular distribution of $w$ around the pillars to emulate the disruption of wave fronts by Marangoni-driven fluid flows (refer to Materials and Methods for more information). The results of this analysis are presented in Figure 12C and Figure S6. Depending on the asymmetry introduced, a varying number of petals emerge following at least one initial cycle of circular waves. Remarkably, even with a set asymmetry of $w$ near the obstacles, the orientation of the petals can vary as the simulation progresses. Finally, to emulate the experimental setup of a tilted Petri dish (Figure S5A-B), we adapted the Tyson-Fife model to include a term that reflects a linear increase in fluid thickness (refer to Materials and Methods). In line with our experimental findings, the waves initially emerge in areas with thinner fluid and then move toward regions with thicker fluid (Figure S5C).
3 Conclusions
@@
In this study, we introduce a periodic array of hydrophilic cylindrical obstacles into the CHD-BZ reaction, which rapidly become synchronized centers of wave activity. The thin layer of chemical solution ascending the obstacle surfaces is essential for enabling these structures to function as wave sources, thus precisely determining their spatial positions (Figure 1C). Although an isolated obstacle can also act as a wave center (Figure 7), an array arrangement effectively prevents the formation of random wave centers elsewhere in the solution. Under covered conditions, where evaporation is minimized, wavefronts maintain stable circular shapes through multiple cycles. In contrast, uncovered conditions—with significant evaporation—induce Marangoni-driven fingering instabilities, causing wavefront breakage and the emergence of striking flower-like patterns (Figure 1F). The number of petals in these patterns scales linearly with obstacle diameter (Figure 4), and there is a minimum diameter required for this instability to manifest (Figure S3). Additionally, control experiments where the weight of PDMS molds was monitored after each five-minute wave recording show negligible weight changes, conclusively ruling out any influence of PDMS swelling on our experimental results (Figure S8).
@@
The behavior of a film climbing the exterior of a vertical or an inclined wall, influenced by the opposing forces of a thermally/chemically driven surface tension gradient and gravity, is extensively explored in the literature through numerical analysis using a thin-film model (see review articles [51, 52, 53] and references therein). In the case of a cylindrical obstacle, these models depend on three parameters: the cylinder radius, the upstream film thickness, and the downstream precursor film thickness. As the cylinder radius approaches infinity, the model converges to the scenario of a Marangoni-driven film ascending a vertical plate [54, 55, 56, 78]. Ref. [57] investigates the stability of the contact line by introducing an infinitesimal two-dimensional perturbation to the traveling wave solution, characterized by sinusoidal variations in the azimuthal direction. Through linear stability analysis, the study determines that the traveling wave (the so called Lax shock) is unstable to long-wave perturbations at the contact line. Given the periodic nature of the contact line in the azimuthal direction, only an integer number of fingers can form along it. The analysis identifies the most unstable mode and a neutrally stable cutoff mode, beyond which all modes are stable. These modes increase linearly with the cylinder radius, indicating that the number of fingers formed along the contact line correlates with the cylinder’s circumference. There exists a critical cylinder radius below which the contact line remains stable, with no fingers forming. Our experimental results provide direct validation of these theoretical predictions. In the future, we aim to expand the numerical analysis from Ref.[57] to include the rounded rectangular obstacle geometry (Figure 8), incorporating an asymmetric evaporation rate that results in an asymmetric surface tension gradient. Additionally, we will integrate the thin film equations from Ref. [57] with the reaction-diffusion equations to fully capture the dynamics of both the fluid flow and chemical components in the thin film surrounding the obstacles.
@@
In our experiments, by tilting the Petri dish, it is possible to selectively control which array of obstacles initially becomes the focal point for wave generation (Figure S5A-B). As reported in Ref.[31] and consistent with our experiments, in a tilted Petri dish, waves originate in the thinner areas of the fluid, where convective flows are negligible, and progress toward the thicker regions, where convective flows can develop due to evaporative cooling. They assumed that these convective flows generate a heterogeneous medium where the refractory tail of a preceding wave annihilates part of the following wavefront, resulting in segmented waves that curl inward to form spirals. Over time, a chaotic wave pattern develops. In another study to understand the role of the fluid thickness in an uncovered Petri dish, Rossi et al.[40] have demonstrated that for fluid layers thinner than 1.0 mm, the BZ system behaves predominantly as a pure reaction-diffusion system, and convective effects are negligible. In layers with an intermediate thickness of 1.0–3.0 mm (the fluid thickness in our experiments was about 1.5 mm), Marangoni instabilities dominate, allowing rippled waves to propagate without breaking, and segmented patterns are not observed. However, for layers thicker than 3.0 mm, buoyancy-driven flows become significant, and wave segmentation processes occur. Our experiment using an uncovered, plasma-treated glass Petri dish with a fluid thickness of 2 mm also revealed rippled wavefronts and Marangoni-driven mosaic patterns (Figure S7A and Video 23) [42, 43, 44, 45]. We notice that at the boundary of the Petri dish, we observed a Marangoni-driven fingering instability, similar to the phenomena we see near the hydrophilic obstacles.
@@
Finally, in our experiments with an evaporating CHD-BZ droplet, we observed the formation of spots at the rim. To elucidate the underlying mechanism, identify the more volatile components, and determine which chemical elements are segregated at the boundary, additional comprehensive experiments are necessary. Tracer beads added to the CHD-BZ solution in our evaporating droplet experiment reveal a pinned contact line and the coffee-stain effect, further suggesting the presence of compensatory capillary flows. Interestingly, after approximately one hour, we observe a complex secondary instability at the droplet boundary, which warrants further investigation. Ongoing studies in our laboratory are focused on exploring this phenomenon in greater detail.
In summary, our study on the spatio-temporal dynamics of chemical waves interacting with a periodic array of obstacles highlights several potential applications. By demonstrating how these obstacles can synchronize and guide wave propagation, we introduce a strategy for engineering specific wave patterns that could be especially beneficial in advanced microfluidic devices. We further propose that these guided patterns can be harnessed not only for exploring fundamental reaction–diffusion dynamics, but also for practical applications such as chemical computing, where controlled wave interactions serve as the basis for logic operations and information processing [81]. Moreover, our findings on manipulating Marangoni-driven flows around hydrophilic obstacles to shape wave patterns can be extended to colloidal droplet evaporation. By employing similar obstacle-based or surface-modification techniques, one could direct evaporative flows and steer colloidal self-assembly, opening new possibilities for creating uniform, tunable deposition patterns in printing, photonics, and related applications [82]. Finally, by adjusting parameters like obstacle geometry or evaporation conditions, one could similarly guide chemical wave propagation within gels, thereby influencing their mechanical oscillations. This capability paves the way for designing advanced biomimetic actuators or “smart” materials where desired wave-driven shape changes (or gel motions) can be precisely programmed or directed [83]. Overall, our results provide a foundation for further manipulation of chemical wave behaviors, offering exciting opportunities for technology and manufacturing [21, 84].

### v4

@@ 2.2 Estimation of dimensionless numbers and timescale comparison
In our experiments, the chemical fingers forming on the pillars or on the walls of the PDMS/PMMA setups exhibit a wavelength of approximately $\lambda\approx 2~mm$, see Figure 4F, inset. We hypothesize that the fluid fingers have a similar wavelength. Assuming the climbing fluid depth is in the range $h=20-100~\mu$[81, 73], we calculate the Marangoni stress $\tau$ from the balance with gravity, $\tau=2\rho gh/3$ [70], which represents the balance between the competing Marangoni stress and gravitational forces. This calculation yields a surface stress range of $\tau=0.13-0.67$ $N/m^{2}$. Evaluating the corresponding solutal Marangoni number for $\tau=0.67$ $N/m^{2}$ and $h=20~\mu m$ yields:
$\mathscr{M}_{S}=\frac{\tau h^{2}}{\mu D}\sim\frac{0.67Nm^{-2}\times(20\times 10^{-6})^{2}m^{2}}{10^{-3}Nsm^{-2}\times 1.2\times 10^{-9}m^{2}s^{-1}}\sim 250,$
where $\mu$ is the viscosity of the CHD-BZ solution and $D$ is molecular diffusivity of chemical components. The corresponding thermal Marangoni number, defined as:
@@
is two orders of magnitude smaller due to the ratio of molecular diffusivity to thermal diffusivity, given by the inverse Lewis number
[…]
Indeed, in general, for evaporating multicomponent liquids with different volatilities of their components, solutal Marangoni effects are more relevant than thermal ones [82, 83]. We now proceed by considering a dimensional analysis similar to Ref. [73] that examines the interaction between a deformable free surface and the steady-state concentration profile, which is governed by diffusion, evaporation, and convective transport along the pillars. Evaporation reduces the concentration of the more volatile chemical component with lower surface tension near the surface; however, under the assumption of a quasi-steady state, fresh fluid from the reservoir continuously replenishes the depleted region, establishing a stable vertical concentration gradient. We note that given the complexity of the CHD-BZ reaction, where multiple chemical components are continuously produced and consumed, the validity of this assumption requires further extensive investigation in future studies. If the free surface experiences an upward perturbation, the concentration profile must adjust before the interface returns to its original shape. If the concentration realigns more rapidly than the interface relaxes, the perturbed region develops a lower concentration than its surroundings, generating a surface tension gradient that draws fluid into the bump and amplifies the perturbation. Similarly, depressions in the free surface create local surface tension minima, pulling fluid away and reinforcing the deformation. This feedback loop sustains convection and enhances surface fluctuations, making the film susceptible to instability driven by Marangoni forces. The system’s stability depends on the relative timescales of diffusion and interface relaxation. If the diffusive timescale is significantly shorter than the relaxation time, the system is unstable, i.e., $T_{\mathrm{diff}}/T_{\mathrm{relax}}\ll 1.$ The relaxation time $T_{\mathrm{relax}}$ is estimated by balancing pressure gradients with viscous stresses,
@@ 2.3 Numerical Simulations
Comprehensive numerical modeling of the CHD–BZ reaction—including its full chemical kinetics—is beyond the scope of this manuscript and will be pursued in future work. To capture the Marangoni-driven hydrodynamic instability qualitatively, we instead employ a simplified two-component fluid mixture exhibiting Marangoni instabilities – e.g., a binary ethanol–water mixture, in which ethanol is both more volatile and has lower surface tension than water. We then reproduce the experimental geometry by placing this mixture around a pillar (Fig. 12) and considering diffusion-limited evaporation of both components. The contact line remains pinned to the pillar walls, and a symmetry boundary condition is imposed on the right. Our evaporation model follows the approach of Section 2.1 in Ref. [83], suitably adapted for this configuration (see Method Section 4).
Figure 12: (A) Geometry of the problem setup based on the experiments at time tt=599 s (see Fig. 6E). The liquid layer (magenta) is pinned to the pillar walls, forming an angle $\theta$ = 52 with the pillar walls and it is at a height of $2.5\text{\,}\mathrm{mm}$ at the walls. The liquid evaporates through diffusion of vapor in the gas phase, as schematically represented by the green arrows. Note the longer arrows close to the contact line, representing the larger evaporation flux close to the contact line singularity. (B) Schematics of the Marangoni instability phenomena in a box with periodic side boundary and an evaporating top surface, where only one of the two components evaporates. When the evaporating component has a lower surface tension, the interface increases its concentration on higher surface tension fluid. May a perturbation locally reduce the surface tension somewhere in the interface, it will onset a Marangoni flow, which in turn draws fluid away from the perturbed region (yellow arrows in second panel). Due to continuity, the displaced fluid must be replenished, which brings fluid from the bulk towards the perturbed region (green arrow in second panel). This draws lower surface tension fluid to the perturbed region, further decreasing the surface tension and creating a positive feedback loop that destabilizes the system – leading to the so-called Marangoni instability.
First, we compute the axisymmetric quasi-stationary base state as a function of the Marangoni number, $\mathrm{Ma}$. Next, we carry out a linear stability analysis with respect to azimuthal perturbations of wavenumber $m$. Our results reveal that the radial width of the liquid layer strongly influences stability: wider layers support longer-wavelength modes. As illustrated schematically in Figure 12B (see also Figure 1 of Ref. [83]), a local drop in surface tension drives Marangoni flow away from the perturbed region. The resulting pressure gradient pulls fluid – enriched in the low-surface tension component – from the bulk back into that region, creating a positive feedback loop that can trigger the instability. The resulting positive feedback can destabilize the system, leading to a Marangoni instability. A wider liquid layer reduces flow restrictions, thereby allowing longer-wavelength perturbations to develop.
To specifically assess the role of the Marangoni number in this confined geometry and avoid varying geometry effects, the liquid width is fixed at 3.5 mm. We then vary $\mathrm{Ma}$ and analyze the stability of the base compositional state with respect to perturbations of different azimuthal wavenumbers. Figure 13 shows the real eigenvalues of the system as functions of both mm and $\mathrm{Ma}$. The results reveal that the system remains stable for low $\mathrm{Ma}$ values ($\mathrm{Ma}\lessapprox 100$) but becomes unstable as $\mathrm{Ma}$ increases. At low $\mathrm{Ma}$, the dominant instabilities arise from perturbations with lower $m$, while higher $\mathrm{Ma}$ shifts the instability toward larger $m$.
Figure 13: Dispersion relation of the system for different $\mathrm{Ma}$ values. The real part of the nondimensionalized eigenvalues is shown as a function of the azimuthal wavenumber $m$ for $\mathrm{Ma}$ values of 100, 2500, 5000, 7500, and 10000. The $\mathrm{Re}(\lambda)=0^{-}$ line indicates the stability threshold, where the real part of the eigenvalue becomes positive, indicating instability.
To visualize a fully three‐dimensional compositional field, we can plot
$y(r,z,\psi)\;=\;y^{0}(r,z)\;+\;\epsilon\,y^{m}(r,z)\,e^{im\psi}\quad+\mathrm{c.c.},$
where $\epsilon$ is a small perturbation amplitude, $y^{0}$ is the axisymmetric base state, and $y^{m}$ is the eigenfunction for azimuthal mode $m$, see Figure 14 showing the full flow (panel A) and the mode $m=5$ (panels B-D). This representation assumes a single dominant eigenmode – which may not strictly hold here – but serves to qualitatively illustrate the emerging instability pattern.
Figure 14: (A) Nondimensional axisymmetric base solution at $\mathrm{Ma}=2500$ for liquid’s velocity magnitude (left), liquid’s composition (right) and gas vapor’s concentration. In the liquid-gas interface, the arrows shown represent the evaporation rate. (B) Eigenfunction with azimuthal mode $m=5$ for the liquid’s $(r,z)$-velocity magnitude (left) and liquid’s composition (right). (C) Top view of the three-dimensional expansion $y(r,z,\psi)=y^{0}(r,z)+\epsilon y^{m}(r,z)e^{im\psi}+\mathrm{c.c.}$ for the composition. (D) Isometric view of the same expansion.
3 Conclusions and outlook
@@
In this study, we introduce a periodic array of hydrophilic cylindrical obstacles into the CHD-BZ reaction, which rapidly become synchronized centers of wave activity. The thin layer of chemical solution ascending the obstacle surfaces is essential for enabling these structures to function as wave sources, thus precisely determining their spatial positions (Figure 1C). Although an isolated obstacle can also act as a wave center (Figure 7), an array arrangement effectively prevents the formation of random wave centers elsewhere in the solution. Under covered conditions, where evaporation is minimized, wavefronts maintain stable circular shapes through multiple cycles. In contrast, uncovered conditions—with significant evaporation—induce Marangoni-driven instabilities, causing wavefront breakage and the emergence of striking flower-like patterns (Figure 1F). The number of petals in these patterns scales linearly with obstacle diameter (Figure 4), and there is a minimum diameter required for this instability to manifest (Figure S3). Additionally, control experiments where the weight of PDMS molds was monitored after each five-minute wave recording show negligible weight changes, conclusively ruling out any influence of PDMS swelling on our experimental results (Figure S7).
We performed numerical simulations of a simplified binary fluid mixture exhibiting Marangoni instabilities (for example, an ethanol–water system) to investigate the evaporation-driven hydrodynamic instability. Convective flows in this system arise from selective evaporation of one component (ethanol) and the resulting surface-tension gradients at the air–liquid interface. In these qualitative simulations—where we assume that the contact line of the fluid mixture is pinned around a single pillar—we show that, in principle, such a simplified ethanol–water mixture can also exhibit a “compositional flower” pattern surrounding the pillar. We stress that these simulations are intended only for qualitative insight and do not quantitatively reproduce the experiments, which involve manyadditional complexities. Moreover, they do not capture thin-film dynamics or fingering instabilities that could emerge once the contact line moves; those phenomena will be addressed in future work.
@@
The behavior of a film climbing the exterior of a vertical or an inclined wall, influenced by the opposing forces of a thermally/chemically driven surface tension gradient and gravity, is extensively explored in the literature through numerical analysis using a thin-film model (see review articles [51, 52, 53] and references therein). In the case of a cylindrical obstacle, these models depend on three parameters: the cylinder radius, the upstream film thickness, and the downstream precursor film thickness. As the cylinder radius approaches infinity, the model converges to the scenario of a Marangoni-driven film ascending a vertical plate [54, 55, 56, 73]. Ref. [57] investigates the stability of the contact line by introducing an infinitesimal two-dimensional perturbation to the traveling wave solution, characterized by sinusoidal variations in the azimuthal direction. Through linear stability analysis, the study determines that the traveling wave (the so called Lax shock) is unstable to long-wavelength perturbations at the contact line. Given the periodic nature of the contact line in the azimuthal direction, only an integer number of fingers can form along it. The analysis identifies the most unstable mode and a neutrally stable cutoff mode, beyond which all modes are stable. These modes increase linearly with the cylinder radius, indicating that the number of fingers formed along the contact line correlates with the cylinder’s circumference. There exists a critical cylinder radius below which the contact line remains stable, with no fingers forming. Our experimental results provide direct validation of these theoretical predictions. In future work, we will extend the thin-film framework of Smolka et al.[57] to rounded-rectangular obstacles—introducing asymmetric evaporation rates and surface-tension gradients—and couple those thin-film equations with the reaction–diffusion equations[6], to capture the fully coupled fluid–chemical dynamics in the thin film around obstacles of various geometries.
@@
In our experiments, by tilting the Petri dish, it is possible to selectively control which array of obstacles initially becomes the focal point for wave generation (Figure S5A-B). As reported in Ref.[32] and consistent with our experiments, in a tilted Petri dish, waves originate in the thinner areas of the fluid, where convective flows are negligible, and progress toward the thicker regions, where convective flows can develop due to evaporation and the resulting Marangoni forces. The authors of that paper assumed that these convective flows generate a heterogeneous medium where the refractory tail of a preceding wave annihilates part of the following wavefront, resulting in segmented waves that curl inward to form spirals. Over time, a chaotic wave pattern develops. In another study to understand the role of the fluid thickness in an uncovered Petri dish, Rossi et al.[41] have demonstrated that for fluid layers thinner than 1.0 mm, the BZ system behaves predominantly as a pure reaction-diffusion system, and convective effects are negligible. In layers with an intermediate thickness of 1.0–3.0 mm (the fluid thickness in our experiments was about 1.5 mm), Marangoni instabilities dominate, allowing rippled waves to propagate without breaking, and segmented patterns are not observed. However, for layers thicker than 3.0 mm, buoyancy-driven flows become significant, and wave segmentation processes occur. This resembles the competition between buoyancy-driven flow and solutal Marangoni in evaporating binary droplets [80]. Our experiment using an uncovered, plasma-treated glass Petri dish with a fluid thickness of 2 mm also revealed rippled wavefronts and Marangoni-driven mosaic patterns (Figure S7A and Video 23) [43, 44, 45, 46]. We notice that at the boundary of the Petri dish, we observed a Marangoni-driven fingering instability, similar to the phenomena we see near the hydrophilic obstacles.
@@
Finally, in our experiments with an evaporating CHD-BZ droplet, we observed the formation of spots at the rim. To elucidate the underlying mechanism, identify the more volatile components, and determine which chemical elements are segregated at the boundary, additional comprehensive experiments are necessary. Tracer beads added to the CHD-BZ solution in our evaporating droplet experiment reveal a pinned contact line and the coffee-stain effect, further suggesting the presence of compensatory capillary flows. Interestingly, after approximately one hour, we observed a complex secondary instability at the droplet boundary, which warrants further investigation. Ongoing studies in our laboratory are focused on exploring this phenomenon in more detail.
Our study on the spatio-temporal dynamics of chemical waves interacting with a periodic array of obstacles highlights several potential applications. By demonstrating how these obstacles can synchronize and guide wave propagation, we introduce a strategy for engineering specific wave patterns that could be especially beneficial in advanced microfluidic devices. We further propose that these guided patterns can be harnessed not only for exploring fundamental reaction–diffusion dynamics, but also for practical applications such as chemical computing, where controlled wave interactions serve as the basis for logic operations and information processing [85]. Moreover, our findings on manipulating Marangoni-driven flows around hydrophilic obstacles to shape wave patterns can be extended to colloidal droplet evaporation. By employing similar obstacle-based or surface-modification techniques, one could direct evaporative flows and steer colloidal self-assembly, opening new possibilities for creating uniform, tunable deposition patterns in printing, photonics, and related applications [86]. Finally, by adjusting parameters like obstacle geometry or evaporation conditions, one could similarly guide chemical wave propagation within gels, thereby influencing their mechanical oscillations. This capability paves the way for designing advanced biomimetic actuators or “smart” materials where desired wave-driven shape changes (or gel motions) can be precisely programmed or directed [87]. Overall, our results provide a foundation for further manipulation of chemical wave behaviors, offering exciting opportunities for technology and manufacturing [21, 88].